OpenStudy (anonymous):

Expand the expression: (y^2-1)^5

OpenStudy (zepp):

Use the binomial theorem to solve this :)

OpenStudy (zepp):

Let's say \(y^2\) is \(a\); and 1 is \(b\).

OpenStudy (zepp):

\(\large (a-b)^5=a^5-5a^4b+10a^3b^2-10a^2b^3+5ab^4-b^5\)

OpenStudy (zepp):

Plug in what you've replaced a and b with \(\large ((y^2)-(1))^5=(y^2)^5-5(y^2)^4(1)+10(y^2)^3(1)^2-10(y^2)^2(1)^3+5(y^2)(1)^4-(1)^5\)

OpenStudy (zepp):

Let's break it into two line, since's not visible. \(\large ((y^2)-(1))^5=\\(y^2)^5-5(y^2)^4(1)+10(y^2)^3(1)^2-10(y^2)^2(1)^3+5(y^2)(1)^4-(1)^5\)

OpenStudy (anonymous):

Thanks!

OpenStudy (zepp):

And simplify! :) You are welcome :)